OpenStudy (anonymous):

An online company sells handmade samurai katana swords. The website costs $ 475 a month to maintain. Each katana costs $ 200 to make, and they sell each katana for $ 755. Create a linear model in the form y = m x + b where x is the number of swords sold per month and y is the net monthly profit. Using this model, find the number of swords that would need to be sold per month to have a monthly profit of $ 4370.

OpenStudy (anonymous):

y=0-485 (the amount made selling 0 swords)+755-200 (the amount made per sword)x y=-475+555x Then... having a profit of $4370 4370=-475+555x 4845/555=9 9=x 9 swords need to be sold to make a profit of $4370.